# Task-driven Image Fusion with Learnable Fusion Loss

### v1

@@ 3.7 Theoretical Analysis
Here, $w_{a},w_{b}\in\mathbb{R}^{H\times W}$, $I_{a},I_{b}\in\mathbb{R}^{H\times W}$, and $\odot$ denotes the element-wise multiplication operation. Let $\Omega^{{}^{\prime}}$ be the set $\{\theta_{\mathcal{F^{\prime}}},\theta_{\mathcal{T^{\prime}}}\}$; this leads to the following expression:
@@
Here, $\mathbf{G}$ denotes the inner product between two gradients: the first one is derived from the task loss using a meta-testing set, and the second is calculated from the fusion loss based on a meta-training set. Consequently, the optimization of the module $\theta_{\mathcal{G}}$ is driven by the task loss, with the objective of preserving task-specific information throughout the fusion process.
Algorithm 1 TDFusion Training Algorithm
1:Training set $\{I_{a}^{ftr},I_{b}^{ftr}\}$ with size N.
2:Thoroughly trained $\theta_{\mathcal{F}}$, $\theta_{\mathcal{T}}$, $\theta_{\mathcal{G}}$.
4:for $epoch=1$ to L do
5:     Sample $\{I_{a}^{mtr},I_{b}^{mtr}\}$ and $\{I_{a}^{mts},I_{b}^{mts}\}$.
6:     for $step=1$ to M do
7:   % Inner update: apply $\mathcal{G}$.
8:         Sample $(I_{a}^{mtr},I_{b}^{mtr})$ and get $(I_{f}^{mtr},y^{mtr})$.
10:   % Outer update: optimize $\mathcal{G}$.
11:         Sample $(I_{a}^{mts},I_{b}^{mts})$ and get $((I_{f}^{mts},y^{mts})$.
12:         Update $\theta_{\mathcal{G}}$ by Eq. 6.
13:     end for
14:     for $step=1$ to N do
15:   % Fusion update: optimize $\mathcal{F}$ and $\mathcal{T}$.
16:         Sample $(I_{a}^{ftr},I_{b}^{ftr})$ and get $(I_{f}^{ftr},y^{ftr})$.
17:         Update $\theta_{\mathcal{F}}$ and $\theta_{\mathcal{T}}$ by Eq. 8 and Eq. 9.
18:     end for
19:end for
@@ 4.1 Setup
Dataset Split. We use four datasets annotated for downstream tasks, including MSRS [44], FMB [26], M3FD [25], and LLVIP [13]. MSRS contains 1083/361 image pairs for training/testing, and FMB contains 1220/280 image pairs for training/testing. We follow the splits of the original papers for both datasets. M3FD dataset consists of 4200 pairs of images for detection, with 300 pairs designated for fusion evaluation. The 4200 detection images are split into 3150 for training and 1050 for testing, ensuring that the 300 pairs for fusion evaluation are included within the detection test set. These 300 fusion images are then employed to assess the fusion performance. The original LLVIP dataset contains 12025/3463 image pairs as training/test set. Due to its large size, we select every 10th image to form our training and test sets, resulting in 1203/347 image pairs for training and testing. Our splits for M3FD and LLVIP will be publicly available.
@@ 4.2 Fusion Experiments
EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q_{abf}$uparrow SSIM $\uparrow$ EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q_{abf}$uparrow SSIM $\uparrow$
@@
EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q_{abf}$uparrow SSIM $\uparrow$ EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q_{abf}$uparrow SSIM $\uparrow$
@@
Fig. 2 presents a visual comparison of different methods. The fused images generated by TDFusion excel in preserving key details, achieving balanced brightness, and maintaining clear object contours. It effectively preserves the target features from the infrared images and the background details from the visible images, resulting in fused images that are more natural and exhibit greater clarity across different environments. These results highlight the advantages of TDFusion in detail preservation and visual performance. More results are available in the supplementary material.
Tab. 1 presents the quantitative comparison of fusion over four datasets. TDFusion outperforms other methods across most metrics. This suggests that TDFusion not only enhances image details but also provides consistent fusion results across various scenarios. Compared to other methods, TDFusion shows superior adaptability and robustness, particularly in handling diverse image characteristics and challenging fusion tasks.
@@ 4.3 Downstream Applications
In this section, for a fair comparison, we use SegFormer [5] and YOLOv8 [46] as backbones and retrain the task networks for each fusion method over 300 epochs to evaluate their adaptability to downstream tasks. Fig. 3 and Fig. 4 present the visual comparisons of semantic segmentation and object detection, respectively. TDFusion outperforms in image detail retention, edge clarity, and object recognition, effectively identifying and segmenting objects. For semantic segmentation, the generated maps clearly distinguish different class regions, closely matching the ground truth. In object detection, the fused images exhibit more precise boundary localization for key objects. This indicates that TDFusion maintains a better balance between fine details and overall context. More results can be found in the supplementary material.
Tab. 2 shows the performance comparison across different methods in semantic segmentation and object detection. TDFusion outperforms other methods on most metrics, particularly in terms of key indicators such as mIoU and mAP. This demonstrates that TDFusion effectively enhances fused image details. It also improves the downstream task performance, especially in terms of accuracy and robustness in complex scenarios. Class-wise results are provided in the supplementary material.
@@ 4.4 Task-driven Learnable Loss
Figure 5: Visual comparison of learnable loss for different downstream tasks.
Our framework incorporates a learnable fusion loss that models the preferences of downstream tasks for information from source images. The models trained on FMB dataset and LLVIP dataset labeled as $SS$ and $OD$, are evaluated on MSRS dataset to simulate performance in unknown scenes, as shown in Fig. 5. The results demonstrate that the fusion model adaptively selects information from infrared and visible images to satisfy task requirements. In semantic segmentation, the model combines scene structure and texture, prioritizing boundaries. This improves segmentation under varying lighting conditions. Fusion weights $w^{ss}_{a}$ and $w^{ss}_{b}$ indicate a preference for visible light details and infrared advantages in low-light conditions. While in object detection, the model focuses on edge and contrast information, especially for targets like pedestrians and vehicles. Higher fusion weights are assigned to bright regions in infrared images, enhancing target detection in low-light conditions. Fusion weights $w^{od}_{a}$ and $w^{od}_{b}$ reflect this preference. Comparison of fusion losses across tasks reveals distinct differences, especially in highlighted regions, confirming that the model adapts to task-specific requirements by selecting the most relevant information from multimodal images. More results are provided in the supplementary material.
@@
4.5 Ablation Studies
To thoroughly evaluate the performance of our proposed algorithm, we conducted a series of ablation experiments on FMB dataset, and the detailed results are shown in Tab. 3. In Experiment I, we excluded the learnable fusion loss by fixing $w_{a}$ and $w_{b}$ to $1/2$. In Experiment II, we omitted the gradient loss from the loss function. In Experiment III, we also allowed the fusion module parameters to be jointly optimized by both the task loss $\mathcal{L}_{t}$ and fusion loss $\mathcal{L}_{f}$. In Experiment IV, we excluded the fusion module’s dedicated learning phase and updated it during the outer update of the loss module, which tests the impact of the fusion module’s training schedule. In Experiment V, we replaced our fusion method with a discriminative approach. The performance decline observed across different configurations confirms the rationality and effectiveness of our proposed method.

### v2

@@ 3.7 Theoretical Analysis
Here, $w_{a},w_{b}\in\mathbb{R}^{H\times W}$, $I_{a},I_{b}\in\mathbb{R}^{H\times W}$, and $\odot$ denotes the element-wise multiplication operation. Let $\Omega^{{}^{\prime}}$ be the set $\{\theta_{\mathcal{F^{\prime}}},\theta_{\mathcal{T^{\prime}}}\}$, this leads to the following expression:
@@
Here, $\mathbf{G}$ denotes the inner product between two gradients: (a) the first one is derived from the task loss using a meta-testing set, and (b) the second is calculated from the fusion loss based on a meta-training set. Consequently, the optimization of the module $\theta_{\mathcal{G}}$ is driven by the task loss, with the objective of preserving task-specific information throughout the fusion process.
@@ 4.1 Setup
Dataset Split. We use four datasets annotated for downstream tasks, including MSRS [58], FMB [34], M3FD [32], and LLVIP [17]. MSRS contains 1083/361 image pairs for training/test, and FMB contains 1220/280 pairs for training/test. We follow the splits of the original papers for both datasets. M3FD dataset consists of 4200 pairs of images for detection, with 300 pairs designated for fusion evaluation. The 4200 detection images are split into 3150 for training and 1050 for testing, ensuring that the 300 pairs for fusion evaluation are included within the detection test set. These 300 fusion images are then employed to assess the fusion performance. The original LLVIP dataset contains 12025/3463 image pairs as training/test set. Due to its large size, we select every 10th image to form our training and test sets, resulting in 1203/347 image pairs for training and testing. Our splits for M3FD and LLVIP will be available.
@@ 4.2 Fusion Experiments
EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q^{AB/F}$$\uparrow$ SSIM $\uparrow$ EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q^{AB/F}$$\uparrow$ SSIM $\uparrow$
@@
EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q^{AB/F}$$\uparrow$ SSIM $\uparrow$ EN $\uparrow$ SF $\uparrow$ SCD $\uparrow$ VIF $\uparrow$ $Q^{AB/F}$$\uparrow$ SSIM $\uparrow$
@@
Fig. 2 presents a visual comparison of different methods. The fused images generated by TDFusion excel in preserving discriminative details, achieving balanced brightness, and maintaining clear object contours. It effectively preserves the target features from the infrared images and the background details from the visible images, resulting in fused images that are more natural and exhibit greater clarity across different environments. These results highlight the advantages of TDFusion in detail preservation and visual performance. More results are available in the supplementary material. Tab. 1 presents the quantitative comparison of fusion over four datasets. TDFusion outperforms other methods across most metrics. This suggests that TDFusion not only enhances image details but also provides consistent fusion results across various scenarios. Compared to other methods, TDFusion shows superior adaptability and robustness, particularly in handling diverse image characteristics and challenging fusion tasks.
@@ 4.3 Downstream Applications
This section validates the adaptability of fusion methods to downstream tasks. For a fair comparison, we adopt SegFormer [9] and YOLOv8 [19] as backbones and retrain the task networks for each fusion method over 300 epochs to evaluate their adaptability to semantic segmentation and object detection. Fig. 3 and Fig. 4 present the visual comparisons of semantic segmentation and object detection, respectively. TDFusion outperforms in image detail retention, edge clarity, and object recognition, effectively identifying and segmenting objects. For semantic segmentation, the generated maps clearly distinguish different class regions, closely matching the ground truth. In object detection, the fused images exhibit more precise boundary localization for salient objects. This indicates that TDFusion maintains a better balance between fine details and overall context. More results can be found in the supplementary material.
@@
Tab. 2 shows the performance comparison across different methods in semantic segmentation and object detection. TDFusion outperforms other methods on most metrics, particularly in mIoU and mAP. This demonstrates that TDFusion effectively enhances the quality of fused images. It also improves the downstream task performance, especially in terms of accuracy and robustness in complex scenarios. Class-wise results are provided in the supplementary material.
@@ 4.4 Task-driven Learnable Loss
Figure 5: Visualisation of learnable loss for downstream tasks.
Our framework incorporates a learnable fusion loss that models the preferences of downstream tasks for information from source images. The models trained on FMB dataset and LLVIP dataset labeled as $SS$ and $OD$, are evaluated on MSRS dataset to simulate performance in unknown scenes, as shown in Fig. 5. The results demonstrate that the fusion model adaptively selects information from infrared and visible images to satisfy task requirements. In semantic segmentation, the model combines scene structure and texture, prioritizing boundaries. This improves segmentation under varying lighting conditions. Fusion weights $\{w^{SS}_{a},w^{SS}_{b}\}$ indicate a preference for visible details and infrared advantages in low-light conditions. In object detection, the model focuses on edge and contrast information, especially for instances like pedestrians and vehicles. Higher fusion weights are assigned to bright regions in infrared images, enhancing target detection in low-light conditions. Fusion weights $\{w^{OD}_{a},w^{OD}_{b}\}$ reflect this preference. Comparison of fusion losses across tasks reveals distinct differences, especially in highlighted regions, confirming that the model adapts to task-specific requirements by selecting the most relevant information from multimodal images. More results are provided in the supplementary material.
4.5 Ablation Studies
To thoroughly evaluate the performance of our proposed algorithm, we conduct a series of ablation experiments on FMB dataset, and the detailed results are shown in Tab. 3.
@@ 4.5 Ablation Studies
In Exp. I, we exclude the learnable fusion loss by fixing $w_{a}$ and $w_{b}$ to $1/2$. In Exp. II, we omit the gradient loss from the loss function. In Exp. III, we also allow the fusion module parameters to be jointly optimized by both the task loss $\mathcal{L}_{t}$ and fusion loss $\mathcal{L}_{f}$. In Exp. IV, we exclude the fusion module’s dedicated learning phase and update it during the outer update of the loss module, which tests the impact of the fusion module’s training schedule. In Exp. V, we replace our fusion method with a discriminative approach. The performance decline observed across different configurations confirms the rationality and effectiveness of our proposed method. Visualization and more analysis are provided in the supplementary material.
@@ 5 Conclusion
Acknowledgement
This work has been supported by the National Natural Science Foundation of China under Grant 12201497 and 12371512.